# Coderivative-Based Newton Methods in Structured Nonconvex and Nonsmooth Optimization

### v2

@@ 7.1 Nonconvex Least Square Regression Models
[Proof.] To verify (i), observe that since $g$ is bounded from below by 00, it follows from [55, Exercise 1.24] that $g$ is prox-bounded with threshold $\lambda_{g}=\infty$. By the calculation in (7.4), $\nabla f$ is Lipschitz continuous on ${\rm I\!R}^{n}$ with modulus $L_{f}=\lambda_{\text{\rm max}}(A^{*}A)+2\mu_{2}$ expressed via the largest eigenvalue of $A^{*}A$. Therefore, Assumption 1 is guaranteed for (7.1). Applying now Theorem 6.1 tells us that any accumulation point of $\{x^{k}\}$ is a critical point of (7.1), which therefore justifies (i).
To prove (ii), we know from (i) that the accumulation point $\bar{x}$ is critical for (7.1), which implies that $0\in\partial\varphi(\bar{x})=\nabla f(\bar{x})+\partial g(\bar{x})$, i.e., $-\nabla f(\bar{x})\in\partial g(\bar{x})$. It follows from Theorem 7.3 and the formula $\nabla f(\bar{x})=A^{*}(A\bar{x}-b)+2\mu_{2}\bar{x}$ that condition (7.19) yields the subdifferential continuity of $g$ at $\bar{x}$ for $-\nabla f(\bar{x})$. The prox-regularity of $g$ at $\bar{x}$ for $-\nabla f(\bar{x})$ is a consequence of the variational convexity of $g$ at $\bar{x}$ for $-\nabla f(\bar{x})$.
Finally, we verify the tilt stability of $\bar{x}$ if either $\mu_{2}>0$, or (7.20) holds. To proceed for $\mu_{2}>0$, observe from Corollary 7.2 that $\partial^{2}\varphi(\bar{x},0)$ is positive-definite, and thus $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ by [49, Theorem 1.3]. In the case where (7.20) holds, for any $w\in{\rm I\!R}^{n}$ such that $0\in\partial^{2}\varphi(\bar{x},0)(w)$, we get
$-\nabla^{2}f(\bar{x})\in\partial^{2}g(\bar{x},-\nabla f(\bar{x}))(w),$
@@
and so $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ by [49, Theorem 1.3]. The graph of $\partial g$ is the union of finitely many closed convex sets, which guarantees by [13] that $\partial g$ is semismooth on its graph. All of this allows us to deduce from Theorem 6.1 the $Q$-superlinear convergence of the iterative sequence $\{x^{k}\}$ generated by Algorithm 4 to the solution $\bar{x}$ and thus completes the proof of the theorem.
@@
In the conducted numerical experiments, we considered different scenarios obtained by changing the regularization terms $\mu_{0}$, $\mu_{2}$ and the size of $A$ keeping a constant column-to-row ratio of $n/m=5$ for the matrix $A$, which is generated randomly with i.i.d. (identically and independent distributed) standard Gaussian entries, where $b$ is generated randomly with values of components from 00 to $1$. For all the algorithms and problems, we used $x^{0}=0$ as the starting iterate. To evaluate the performance of our proposed method, GCNM, we run GCNM, GSSN, and ZeroFPR until the condition $\|x^{k}-\widehat{x}^{k}\|\leq 10^{-6}$ is satisfied.
In both of these experiments, we consider the following two cases. In the first case where $\mu_{2}=0.01>0$, generalized Newton directions of GCNM and GSSN always exist, and the superlinear convergence of both algorithms is guaranteed. In the second case where $\mu_{2}=0$, this guarantee does not longer hold, and superlinear convergence may not be achieved. More detailed information for the testing data used in the experiments is presented in Table 1, where “TN” stands for test number. The result for the tests is presented in Table 2, where CPU time stands for the time needed (in seconds) to achieve the prescribed accuracy of approximate solutions.
@@
TN CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$
1 0.019 0.00153 1.2E-17 0.065 0.00181 3.9E-18 6.099 0.00323 2.1E-07
2 0.003 0.00085 2.3E-17 0.011 0.00087 6.7E-18 2.792 0.00101 7.9E-07
3 0.021 0.00130 2.6E-17 0.043 0.00161 1.7E-18 6.443 0.00958 6.0E-07
4 0.013 0.00057 2.8E-17 0.039 0.00059 1.7E-17 10.721 0.00069 6.8E-07
5 0.057 0.00091 3.8E-17 0.430 0.00128 9.2E-18 18.711 0.03326 9.0E-07
6 0.049 0.00040 2.7E-17 0.239 0.00041 2.0E-17 41.580 0.00087 8.6E-07
7 0.438 0.00104 6.9E-17 3.057 0.03252 9.2E-18 186.717 0.06447 8.9E-07
8 0.156 0.00027 3.1E-17 0.688 0.00027 2.5E-17 38.567 0.00023 8.3E-07
9 3.992 0.00067 6.0E-17 8.872 0.00159 1.2E-17 408.838 0.00032 9.4E-07
10 1.379 0.00020 6.3E-17 4.037 0.00020 3.7E-17 238.956 0.00018 9.4E-07
@@
TN CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$
1 0.002 5.2E-15 5.4E-15 0.012 2.3E-15 8.5E-18 0.081 1.2E-05 4.9E-07
2 0.003 4.0E-14 1.3E-14 0.107 8.9E-06 1.3E-07 0.215 1.5E-05 8.4E-07
3 0.003 1.5E-14 5.2E-15 0.091 7.2E-02 7.2E-18 0.809 2.8E-05 8.1E-07
4 0.016 1.5E-14 4.4E-15 1.394 8.6E-05 7.8E-07 0.878 2.9E-05 7.5E-07
5 0.028 3.0E-12 1.1E-12 0.365 2.3E-02 5.2E-18 2.051 5.9E-05 8.1E-07
6 0.016 9.9E-14 2.8E-14 1.577 5.6E-05 4.4E-07 1.774 4.7E-05 6.9E-07
7 0.062 1.3E-13 8.9E-14 2.068 2.2E-01 3.1E-17 10.382 1.2E-04 9.1E-07
8 0.145 5.8E-11 1.9E-11 9.959 1.2E-04 4.9E-07 7.403 8.2E-05 9.2E-07
9 0.571 1.9E-11 4.9E-12 14.230 1.2E-01 3.3E-17 41.439 1.7E-04 8.7E-07
10 0.297 6.7E-13 1.5E-13 34.760 2.9E-04 9.0E-07 26.853 1.0E-04 8.2E-07
@@
In many practical applications, the optimization problem (7.1) is employed to find an approximate solution to the underdetermined linear system $Ax=b$, where $A$ is an $m\times n$ matrix with $m<n$. To assess the effectiveness of approximate stationary points obtained from solving (7.1), we evaluate the quality of the solutions by comparing the norm of the residual $\|Ax^{k}-b\|$ and the deviation from the solution $\|x^{k}-\hat{x}^{k}\|$ after $k$ iterations. The results presented in Table 2 consistently demonstrate the superior performance of GCNM over GSSN and ZeroFPR across various evaluation metrics. In the first case where $\mu_{2}=0.01>0$, GCNM ensures the existence of generalized Newton directions and guarantees superlinear convergence while notably outperforms GSSN in CPU time. This significant difference can be attributed to the simpler construction of the Newton direction in GCNM, which contrasts with the approach in GSSN that requires boundedness of the direction and involves computationally expensive matrix operations. Additionally, the line search strategy used in GSSN tends to be more complex and requires more steps compared to the more efficient line search method implemented in Algorithm 3 for GCNM. Furthermore, in terms of solution quality as discussed earlier, GCNM generally yields smaller residuals $\|Ax^{k}-b\|$. For larger problem instances, ZeroFPR produces smaller residuals $\|Ax^{k}-b\|$; however, this algoritm requires higher computational cost and longer runtimes compared to both GCNM and GSSN. It is also worth noting that in many test problems, GSSN produced smaller residuals $\|x^{k}-\hat{x}^{k}\|$ compared to GCNM and ZeroFPR. The second case, where $\mu_{2}=0$ and superlinear convergence is not guaranteed, further demonstrates the robustness of GCNM. This trend is consistently observed across most tests, with GCNM generally requiring substantially less CPU time.
@@ 7.2 Student’s t‑Regression with $\ell^{0}$-Penalty
The proximal mapping of $g$ in (7.22) is given by (7.5). To run Algorithm 4, we need determining the direction sequence $\{d^{k}\}$ in the Newton step, which requires calculating the second-order subdifferential of $g(x)=\mu\|x\|_{0}$ presented in Proposition 7.1. So we can express $d^{k}$ in Algorithm 4 for solving (7.21) via the conditions
$\begin{cases}\big{(}-\hat{v}^{k}-\nabla^{2}f(\hat{x}^{k})d^{k}\big{)}_{i}=0&\text{if}\quad\left(\hat{x}^{k}\right)_{i}\neq 0,\\ \big{(}d^{k}\big{)}_{i}=0&\text{if}\quad\left(\hat{x}^{k}\right)_{i}=0,\end{cases}$ (7.25)
where the vectors $\hat{x}^{k},\hat{v}^{k}$ are given by
$\hat{x}^{k}=\text{\rm Prox}_{\lambda g}(x^{k}-\lambda\nabla f(x^{k})),\quad\hat{v}^{k}=\nabla f(\hat{x}^{k})-\nabla f(x^{k})+\frac{1}{\lambda}(x^{k}-\hat{x}^{k}),$
@@
Consider problem (7.21) with the representations of $f$ and $g$ given in (7.22) and the explicit computations presented above. Then Algorithm 4 either stops after finitely many iterations, or produces a sequence $\{x^{k}\}$ such that if $\bar{x}$ is an accumulation point of this sequence, then $\bar{x}$ is a critical point of (7.21). If in addition $\varphi$ is subdifferentially continuous and variationally strongly convex at $\bar{x}$ for 00, then $\{x^{k}\}$ Q-superlinearly converges to $\bar{x}$.
[Proof.] Observe that since $g$ is bounded from below by 00, it follows from [55, Exercise 1.24] that $g$ is prox-bounded with threshold $\lambda_{g}=\infty$. Moreover, $f$ is smooth with the gradient and Hessian given in (7.23) and (7.24). It follows from (7.23) that $\nabla f$ is Lipschitz with constant $L=2\|A\|_{1}\|A\|_{\infty}$, where
$\|A\|_{1}=\max_{1\leq j\leq n}\sum_{i=1}^{n}|a_{ij}|,\quad\|A\|_{\infty}=\max_{1\leq i\leq n}\sum_{j=1}^{n}|a_{ij}|,\quad A=[a_{ij}]_{i,j=\overline{1,n}}.$
Therefore, Assumption 1 is fulfilled for problem (7.21). Thus it follows from Theorem 6.1 that if $\bar{x}$ is an accumulation point of $\{x^{k}\}$, then $\bar{x}$ is a critical point of (7.21).
It remains to verify the Q-superlinear convergence of $\{x^{k}\}$ provided that $\varphi$ is subdifferentially continuous and variationally strongly convex at $\bar{x}$ for 00. This ensures that $\varphi$ is continuously prox-regular at $\bar{x}$ for 00, and so $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ by Proposition 2.7. As mentioned above, the graph of $\partial g$ is the union of finitely many closed convex sets, which ensures that $\partial g$ is semismooth on its graph. Therefore, the superlinear convergence to $\bar{x}$ of the iterates $\{x^{k}\}$ generated by Algorithm 4 is guaranteed by Theorem 6.1.
@@
n CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$
40 0.006 2.6E-05 2.9E-06 0.051 5.7E-04 3.3E-05 0.095 3.9E-05 9.0E-05
80 0.003 1.8E-05 8.7E-07 0.548 5.2E-03 9.9E-05 0.045 3.7E-05 8.8E-05
160 0.004 1.5E-05 3.4E-07 4.682 1.7E-02 9.9E-05 0.399 2.6E-04 8.0E-05
320 0.105 5.6E-06 8.7E-08 10.851 1.1E-02 9.9E-05 1.353 4.9E-05 6.4E-05
640 0.294 1.3E-07 1.2E-09 42.929 1.7E-02 9.9E-05 1.078 2.8E-05 7.9E-05
1280 0.460 6.1E-07 2.7E-09 241.641 3.3E-02 9.9E-05 5.629 3.3E-05 9.8E-05
2560 2.472 2.4E-07 5.6E-10 1276.129 5.6E-02 9.9E-05 88.735 1.9E-04 7.3E-05
@@
n CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$
40 0.002 1.4E-06 1.6E-07 0.032 8.0E-04 6.9E-05 0.013 2.8E-05 5.5E-05
80 0.004 9.7E-08 5.5E-09 0.106 2.5E-03 9.9E-05 0.019 3.2E-05 7.8E-05
160 0.010 2.0E-07 5.5E-09 3.484 6.5E-03 9.9E-05 0.102 1.8E-05 5.5E-05
320 0.009 1.6E-09 2.9E-11 2.887 6.8E-03 9.9E-05 0.181 3.1E-05 8.9E-05
640 0.053 3.1E-08 2.8E-10 13.172 1.3E-02 9.9E-05 1.360 5.0E-05 7.7E-05
1280 0.510 2.0E-09 9.2E-12 13.604 2.1E-02 8.7E-05 7.435 3.0E-05 8.3E-05
2560 2.646 1.0E-09 1.1E-12 86.497 1.9E-02 4.2E-05 26.892 1.1E-04 6.6E-05
@@
n CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$ CPU time $\|Ax^{k}-b\|$ $\|x^{k}-\hat{x}^{k}\|$
40 0.002 6.1E-08 6.7E-09 0.053 7.6E-04 7.3E-05 0.010 4.3E-05 8.4E-05
80 0.002 1.4E-10 7.9E-12 0.003 7.3E-04 4.2E-05 0.010 3.2E-05 7.3E-05
160 0.013 4.1E-10 1.2E-11 0.506 4.2E-03 9.9E-05 0.028 1.8E-05 6.8E-05
320 0.005 1.1E-14 2.0E-17 0.002 1.4E-15 8.4E-18 0.010 1.4E-15 9.3E-18
640 0.092 2.5E-10 2.3E-12 5.104 5.3E-03 4.4E-05 1.037 1.9E-05 5.8E-05
1280 0.175 1.5E-10 6.9E-13 14.976 4.5E-03 1.9E-05 1.995 2.0E-05 5.6E-05
2560 2.504 2.9E-10 1.2E-13 84.921 3.9E-03 8.7E-06 14.313 1.8E-05 6.6E-05

### v3

@@ 7.1 Nonconvex Least Square Regression Models
[Proof.] To verify (i), observe that since $g$ is bounded from below by 0, it follows from [54, Exercise 1.24] that $g$ is prox-bounded with threshold $\lambda_{g}=\infty$. By the calculation in (7.4), $\nabla f$ is Lipschitz continuous on ${\rm I\!R}^{n}$ with modulus $L_{f}=\lambda_{\text{\rm max}}(A^{*}A)+2\mu_{2}$ expressed via the largest eigenvalue of $A^{*}A$. Therefore, Assumption 1 is guaranteed for (7.1). Applying now Theorem 6.1 tells us that any accumulation point of $\{x^{k}\}$ is a $\lambda$-critical point of (7.1) for all $\lambda\in(0,1/L_{f})$, which therefore justifies (i).
To prove (ii), we know from (i) that the accumulation point $\bar{x}$ is $\lambda$-critical for (7.1), which implies that $0\in\partial\varphi(\bar{x})=\nabla f(\bar{x})+\partial g(\bar{x})$, i.e., $\nabla f(\bar{x})\in\partial g(\bar{x})$. It follows from Theorem 7.3 and the formula $\nabla f(\bar{x})=A^{*}(A\bar{x}-b)+2\mu_{2}\bar{x}$ that condition (7.16) yields the subdifferential continuity of $g$ at $\bar{x}$ for $\nabla f(\bar{x})$. The prox-regularity of $g$ at $\bar{x}$ for $\nabla f(\bar{x})$ is a consequence of the variational convexity of $g$ at $\bar{x}$ for $\nabla f(\bar{x})$.
Finally, we verify the tilt stability of $\bar{x}$ if either $\mu_{2}>0$, or (7.17) holds. To proceed for $\mu_{2}>0$, observe from Corollary 7.2 that $\partial^{2}\varphi(\bar{x},0)$ is positive-definite, and thus $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ by [48, Theorem 1.3]. In the case where (7.17) holds, for any $w\in{\rm I\!R}^{n}$ such that $0\in\partial^{2}\varphi(\bar{x},0)(w)$, we get
$\nabla^{2}f(\bar{x})\in\partial^{2}g(\bar{x},-\nabla f(\bar{x}))(w),$
@@
and so $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ by [48, Theorem 1.3]. The graph of $\partial g$ is the union of finitely many closed convex sets, which guarantees by [13] that $\partial g$ is semismooth on its graph. All of this allows us to deduce from Theorem 6.1 the $Q$-superlinear convergence of the iterative sequence $\{x^{k}\}$ generated by Algorithm 4 to the solution $\bar{x}$ and thus completes the proof of the theorem.
@@
In the conducted numerical experiments, we considered different scenarios obtained by changing the regularization terms $\mu_{0}$, $\mu_{2}$ and the size of $A$ keeping a constant column-to-row ratio of $n/m=5$ for the matrix $A$, which is generated randomly with i.i.d. (identically and independent distributed) standard Gaussian entries, where $b$ is generated randomly with values of components from 0 to $1$. For all the algorithms and problems, we used $x^{0}=0$ as the starting iterate. To evaluate the performance of our proposed method, GCNM, we run GCNM, GSSN, and ZeroFPR until the condition $\|x^{k}-\widehat{x}^{k}\|\leq 10^{-6}$ is satisfied.
In both of these experiments, we consider the following two cases. In the first case where $\mu_{2}=0.01>0$, generalized Newton directions of GCNM and GSSN always exist, and the superlinear convergence of both algorithms is guaranteed. In the second case where $\mu_{2}=0$, this guarantee does not longer hold, and superlinear convergence may not be achieved. More detailed information for the testing data used in the experiments is presented in Table 1, where “TN” stands for test number. The numerical results of the tests are summarized in Table 2. In the table, “time” denotes the computational time in seconds, “iter” denotes the number of iterations required to attain the prescribed accuracy, $\delta_{k}:=\|Ax^{k}-b\|$, $\eta_{k}:=\|x^{k}-\hat{x}^{k}\|$, and “nnz” denotes the number of nonzero components in the computed approximate solution.
@@
TN time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz
1 0.007 3 7.8E-04 2.9E-17 76 0.044 13 1.2E-03 7.2E-18 33 4.243 455 1.2E-02 6.5E-07 85
2 0.003 3 7.7E-04 2.7E-17 96 0.028 12 7.7E-04 2.0E-17 84 1.747 435 1.0E-01 9.7E-07 94
3 0.021 4 5.7E-04 2.5E-17 170 0.139 12 6.1E-04 1.5E-17 126 5.721 988 7.8E-03 7.4E-07 184
4 0.013 3 5.7E-04 2.3E-17 191 0.039 12 5.7E-04 1.8E-17 162 10.721 1196 1.6E-03 3.7E-07 195
5 0.225 6 4.6E-04 3.4E-17 320 0.757 12 4.8E-04 1.9E-17 240 29.601 1517 4.8E-02 9.1E-07 341
6 0.015 3 4.5E-04 3.2E-17 376 0.359 14 4.5E-04 2.7E-17 318 40.462 2234 1.2E-02 7.2E-07 356
7 0.281 7 2.5E-04 3.8E-17 645 1.266 13 2.8E-04 1.8E-17 392 273.805 2839 3.8E-02 9.8E-07 721
8 0.158 4 2.4E-04 3.5E-17 744 2.215 13 2.4E-04 2.9E-17 605 26.510 266 2.4E-04 8.0E-07 765
9 1.168 5 1.9E-04 5.6E-17 1315 10.298 14 2.0E-04 3.0E-17 877 408.838 1562 3.2E-04 9.2E-07 1455
10 0.630 3 1.9E-04 5.4E-17 1517 9.756 15 1.9E-04 4.8E-17 1370 227.216 362 1.9E-04 9.3E-07 1456
@@
TN time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz
1 0.002 3 5.2E-15 5.4E-15 60 0.012 9 2.3E-15 8.5E-18 27 0.081 37 1.2E-05 4.9E-07 68
2 0.003 4 4.0E-14 1.3E-14 79 0.107 22 8.9E-06 1.3E-07 62 0.215 33 1.5E-05 8.4E-07 99
3 0.003 2 1.5E-14 5.2E-15 109 0.091 10 7.2E-02 7.2E-18 37 0.809 42 2.8E-05 8.1E-07 106
4 0.016 2 1.5E-14 4.4E-15 169 1.394 33 8.6E-05 7.8E-07 110 0.878 36 2.9E-05 7.5E-07 184
5 0.028 3 3.0E-12 1.1E-12 228 0.365 10 2.3E-02 5.2E-18 74 2.051 53 5.9E-05 8.1E-07 176
6 0.016 2 9.9E-14 2.8E-14 339 1.577 27 5.6E-05 4.4E-07 231 1.774 41 4.7E-05 6.9E-07 385
7 0.062 2 1.3E-13 8.9E-14 448 2.068 9 2.2E-01 3.1E-17 148 10.382 49 1.2E-04 9.1E-07 428
8 0.145 5 5.8E-11 1.9E-11 641 9.959 36 1.2E-04 4.9E-07 424 7.403 36 8.2E-05 9.2E-07 757
9 0.571 4 1.9E-11 4.9E-12 891 14.230 10 1.2E-01 3.3E-17 300 41.439 53 1.7E-04 8.7E-07 900
10 0.297 7 6.7E-13 1.5E-13 1303 34.760 32 2.9E-04 9.0E-07 936 26.853 37 1.0E-04 8.2E-07 1501
@@
In many practical applications, the optimization problem (7.1) is employed to find an approximate solution to the underdetermined linear system $Ax=b$, where $A$ is an $m\times n$ matrix with $m<n$. To assess the effectiveness of approximate stationary points obtained from solving (7.1), we evaluate the quality of the solutions by comparing the norm of the residual $\delta_{k}=\|Ax^{k}-b\|$ and the deviation from the solution $\eta_{k}=\|x^{k}-\hat{x}^{k}\|$ after $k$ iterations. The results presented in Table 2 consistently demonstrate the superior performance of GCNM over GSSN and ZeroFPR across various evaluation metrics. In the first case where $\mu_{2}=0.01>0$, GCNM ensures the existence of generalized Newton directions and guarantees superlinear convergence while notably outperforms GSSN in time. This significant difference can be attributed to the simpler construction of the Newton direction in GCNM, which contrasts with the approach in GSSN that requires boundedness of the direction and involves computationally expensive matrix operations. Additionally, the line search strategy used in GSSN tends to be more complex and requires more steps compared to the more efficient line search method implemented
in Algorithm 3 for GCNM. From Table 2, we can observe that the solutions obtained by the different methods differ, as expected for a nonconvex problem. The line search strategy used in GSSN generally requires more iterations, whereas GCNM in these experiments appears to take mostly full Newton steps, resulting in fewer iterations. Furthermore, in terms of solution quality as discussed earlier, GCNM generally yields smaller residuals $\|Ax^{k}-b\|$. For larger problem instances, ZeroFPR produces smaller residuals $\|Ax^{k}-b\|$; however, this algorithm requires higher computational cost and longer runtimes compared to both GCNM and GSSN. It is also worth noting that in many test problems, GSSN produced smaller residuals $\|x^{k}-\hat{x}^{k}\|$ compared to GCNM and ZeroFPR. The second case, where $\mu_{2}=0$ and superlinear convergence is not guaranteed, further demonstrates the robustness of GCNM. This trend is consistently observed across most tests, with GCNM generally requiring substantially less time.
@@ 7.2 Student’s t‑Regression with $\ell^{0}$-Penalty
The proximal mapping of $g$ in (7.19) is given by (7.5). To run Algorithm 4, we need to determine the direction sequence $\{d^{k}\}$ in the Newton step, which requires calculating the second-order subdifferential of $g(x)=\mu\|x\|_{0}$ presented in Proposition 7.1. So we can express $d^{k}$ in Algorithm 4 for solving (7.18) via the conditions
$\begin{cases}\big(-\widehat{v}^{k}-\nabla^{2}f(\widehat{x}^{k})d^{k}\big)_{i}=0&\text{if}\quad\left(\widehat{x}^{k}\right)_{i}\neq 0,\\ \big(d^{k}\big)_{i}=0&\text{if}\quad\left(\widehat{x}^{k}\right)_{i}=0,\end{cases}$
where the vectors $\widehat{x}^{k},\widehat{v}^{k}$ are given by
$\widehat{x}^{k}=\text{\rm Prox}_{\lambda g}(x^{k}-\lambda\nabla f(x^{k})),\quad\widehat{v}^{k}=\nabla f(\widehat{x}^{k})-\nabla f(x^{k})+\frac{1}{\lambda}(x^{k}-\widehat{x}^{k}),$
@@
Consider problem (7.18) with the representations of $f$ and $g$ given in (7.19) and the explicit computations presented above. Then Algorithm 4 either stops after finitely many iterations, or produces a sequence $\{x^{k}\}$ such that if $\bar{x}$ is an accumulation point of this sequence, then $\bar{x}$ is a $\lambda$-critical point of (7.18) for all $\lambda\in(0,1/L_{f})$, where $L_{f}=2\|A\|_{1}\|A\|_{\infty}$. If in addition $\varphi$ is subdifferentially continuous and variationally strongly convex at $\bar{x}$ for 0, then the sequence $\{x^{k}\}$ Q-superlinearly converges to $\bar{x}$.
[Proof.] Observe that since $g$ is bounded from below by 0, it follows from [54, Exercise 1.24] that $g$ is prox-bounded with threshold $\lambda_{g}=\infty$. Moreover, $f$ is smooth with the gradient and Hessian are computed in (7.20) and (7.21), respectively. It follows from (7.20) that $\nabla f$ is Lipschitz with constant $L=2\|A\|_{1}\|A\|_{\infty}$. Therefore, Assumption 1 is fulfilled for problem (7.18). Thus it follows from Theorem 6.1 that if $\bar{x}$ is an accumulation point of $\{x^{k}\}$, then $\bar{x}$ is a $\lambda$-critical point of (7.18) for all $\lambda\in(0,1/L_{f})$.
It remains to verify the Q-superlinear convergence of $\{x^{k}\}$ provided that $\varphi$ is subdifferentially continuous and variationally strongly convex at $\bar{x}$ for 0. This ensures that $\varphi$ is continuously prox-regular at $\bar{x}$ for 0, and so $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ by Proposition 2.8. As mentioned above, the graph of $\partial g$ is the union of finitely many closed convex sets, which ensures that $\partial g$ is semismooth on its graph. Therefore, the superlinear convergence to $\bar{x}$ of the iterates $\{x^{k}\}$ generated by Algorithm 4 is guaranteed by Theorem 6.1.
@@
n time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz
40 0.002 2 2.6E-06 2.9E-05 34 0.023 13 5.7E-04 3.3E-05 7 0.026 15 3.9E-05 9.0E-05 38
80 0.003 3 1.8E-05 8.7E-07 51 0.548 349 5.2E-03 9.9E-05 28 0.045 23 3.7E-05 8.8E-05 51
160 0.004 2 1.5E-05 3.4E-07 85 1.707 703 1.7E-02 9.9E-05 32 0.192 62 2.6E-04 8.0E-05 135
320 0.035 4 5.6E-06 8.7E-08 265 8.181 402 1.1E-02 9.9E-05 160 0.238 25 4.9E-05 6.4E-05 311
640 0.132 5 1.3E-07 1.2E-09 580 34.331 426 1.7E-02 9.9E-05 374 1.335 24 2.8E-05 7.9E-05 581
1280 0.460 3 6.1E-07 2.7E-09 1172 241.641 556 3.3E-02 9.9E-05 775 5.629 15 3.3E-05 9.8E-05 1180
2560 2.472 7 2.4E-07 5.6E-10 2377 1276.129 435 5.6E-02 9.9E-05 1660 88.735 36 1.9E-04 7.3E-05 2435
@@
n time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz
40 0.003 2 1.4E-06 1.6E-07 37 0.032 25 8.0E-04 6.9E-05 18 0.013 10 2.8E-05 5.5E-05 37
80 0.004 3 9.7E-08 5.5E-09 73 0.106 68 2.5E-03 9.9E-05 51 0.019 9 3.2E-05 7.8E-05 73
160 0.010 3 2.0E-07 5.5E-09 128 1.139 162 6.5E-03 9.9E-05 59 0.102 16 1.8E-05 5.5E-05 155
320 0.064 5 1.6E-09 2.9E-11 308 3.356 111 6.8E-03 9.9E-05 247 0.480 10 3.1E-05 8.9E-05 310
640 0.034 3 3.1E-08 2.8E-10 620 10.493 98 1.3E-02 9.9E-05 494 0.709 14 5.0E-05 7.7E-05 621
1280 0.510 3 2.0E-09 9.2E-12 1251 13.604 20 2.1E-02 8.7E-05 1067 7.435 17 3.0E-05 8.3E-05 1255
2560 2.646 4 1.0E-09 1.1E-12 2515 86.497 29 1.9E-02 4.2E-05 2194 26.892 11 1.1E-04 6.6E-05 2525
@@
n time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz time iter $\delta_{k}$ $\eta_{k}$ nnz
40 0.002 2 6.1E-08 6.7E-09 38 0.053 26 7.6E-04 7.3E-05 25 0.010 8 4.3E-05 8.4E-05 38
80 0.002 2 1.4E-10 7.9E-12 79 0.003 2 7.3E-04 4.2E-05 79 0.010 5 3.2E-05 7.3E-05 80
160 0.005 3 4.1E-10 1.2E-11 153 0.123 32 4.2E-03 9.9E-05 113 0.028 7 1.8E-05 6.8E-05 153
320 0.005 2 1.1E-14 2.0E-17 320 0.002 2 1.4E-15 8.4E-18 320 0.010 2 1.4E-15 9.3E-18 320
640 0.092 3 2.5E-10 2.3E-12 632 5.104 24 5.3E-03 4.4E-05 561 1.037 7 1.9E-05 5.8E-05 632
1280 0.175 5 1.5E-10 6.9E-13 1273 14.976 23 4.5E-03 1.9E-05 1171 1.995 7 2.0E-05 5.6E-05 1270
2560 2.504 4 2.9E-10 1.2E-13 2548 84.921 28 3.9E-03 8.7E-06 2383 14.313 6 1.8E-05 6.6E-05 2553